Amber and jenny finish counting their money. Amber says if jenny gives her four dollars they will have the same amount. Jenny says if amber gives her four dollars she will have twice what amber has. how much do they have?

Answer by ankor@dixie-net.com(22740)

Let x = A's $ and y = J's $
:
Write an equation for each statement:
:
Amber and jenny finish counting their money. Amber says if jenny gives her four dollars they will have the same amount.
x + 4 = y - 4
Simplify:
x = y -4 - 4
x = (y-8)
;
Jenny says if amber gives her four dollars she will have twice what amber has.
y + 4 = 2(x-4)
y + 4 = 2x - 8
y = 2x - 8 - 4
y = 2x - 12
:
Substitute (y-8) for x, and find y
y - 2(y-8) - 12
y = 2y - 16 - 12
y - 2y = -28
-y = -28
y = +28 dollars is J's original amt
:
x = y - 8
x = 28 - 8
x = 20 dollars is A,s original amt
:
Check solutions by using these amts in the two statements to see if they are true.